two one kg spheres each carry a charge of magnitude 1c.how does FE,the strength of the electric force between the spheres,compare to Fg,the strength of their gravitational attraction?

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

FE=(1/4pi(ε0))(1*1/r^2)=9*10^9/r^2, Fg=6.67*10^(-11)(1*1)/r2, FE/Fg=9*10^9/(6.67*10^(-11)=1.349*1020